Charles wants to build a vegetable garden such that three sides of the garden are fenced and the fourth side of the garden will be the existing back fence. He has 30 feet of fencing available. Find the dimensions of the garden that will produce the maximum enclosed area.
Answer by checkley77(12844)

30/4=7.5 ft. is the depth of the fence.
So we have a 7.5 by 15 by 7.5 rectangle.
7.5*15=112.5 ft^2 is the ans. for the maximum area.
30/3=10 if it were a square garden.
10*10=100 ft^2.